OpenStudy (anonymous):

Rewrite y=3x² - 24x +10 in general form

OpenStudy (mrsmarsh):

Please contact your FLVS teacher for help with this concept. Posting questions to this forum and copying the answers to your work is cheating and a violation of the Academic Integrity policy which you agreed to when you signed up for the course.

OpenStudy (anonymous):

\[ y = (\sqrt{3}x)^2 - 2 \sqrt 3 4\sqrt 3 x + (4\sqrt 3)^2 + (10 - (4\sqrt 3)^2) \\ = (\sqrt 3x - 4\sqrt 3)^2 - 38 \\= 3 (x - 4)^2 - 38 \]

OpenStudy (anonymous):

6. Create your own quadratic equation and demonstrate how it would be solved by graphing, factoring, the quadratic formula, and by completing the square. (5 points)